# FastSLM: Hierarchical Frame Q-Former for Effective Speech Modality Adaptation

### v1

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated human-expert-level capabilities, driving significant interest in their potential for achieving artificial general intelligence (AGI). In particular, there is growing momentum in adapting LLMs to various modalities—including vision, video, and speech—through the development of multimodal LLMs (MLLMs). However, existing speech-language models (SLMs) research has largely overlooked cost-effective adaptation strategies for leveraging LLMs in the speech domain. In this paper, we propose FastSLM, a lightweight yet efficient SLM designed for effective understanding and reasoning over long-form speech. To address the challenge of aligning high-frame speech features with LLM, we introduce the hierarchical frame querying transformer (HFQ-Former), which compresses frame-level speech features while capturing both local and global context. Furthermore, we present a novel three-stage training strategy that enhances generalization across a wide range of speech-related tasks. Experimental results demonstrate that FastSLM achieves competitive performance compared to existing state-of-the-art (SOTA) models, despite operating with significantly lower FLOPs and parameter counts, while representing speech with only 1.67 tokens per second. The source code and model checkpoints are available at https://huggingface.co/okestro-ai-lab/FastSLM.
@@ 1 Introduction
Recently, large language models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; Comanici et al., 2025; Yang et al., 2025) have demonstrated expert-human level performance in various tasks such as code generation, math, and reasoning (Hendrycks et al., 2020; Wang et al., 2024). Accordingly, in order to move toward the ultimate goal of Artificial General Intelligence (AGI), research on multimodal LLMs (MLLMs) that apply various modalities (vision, speech, video) to LLM is actively being conducted (Yin et al., 2024; Lyu et al., 2023). Among them, speech is one of the core means by which artificial intelligence (AI) communicates with users, so speech recognition and understanding are essential components for achieving AGI (Sakshi et al., 2024). Consequently, much research has been conducted on ALMs (Tang et al., 2024; Abouelenin et al., 2025; Chu et al., 2023; 2024; Goel et al., 2025) that adapt the speech modality to LLM.
Many studies have used frame-level features extracted from pre-trained speech encoders such as Whisper (Radford et al., 2023), CLAP (Elizalde et al., 2023), Conformer (Gulati et al., 2020), and Beats (Chen et al., 2023) as inputs to LLM for speech modality adaptation. This approach is highly effective because it leverages frame-level features from encoders trained on large-scale speech data (Tang et al., 2024; Chu et al., 2023; 2024; Goel et al., 2025). However, projecting hundreds to thousands of frame-level features extracted by the encoder into the LLM using only a multi-layer perceptron (MLP) leads to a significant increase in computational complexity during generation. As the key-value (KV) cache of the autoregressive LLM grows, generation latency also increases, making real-time responses impractical (Arif et al., 2025).
Recently, the importance of long-form speech reasoning has been emphasized in tasks such as speech summarization (SSUM) and spoken-query-based question answering (SQQA) (Ghosh et al., 2024; Kang and Roy, 2024). However, these tasks require processing long-form speech input, which significantly increases the computational complexity of LLM. To address this challenge, we investigate a cost-effective design and training strategy for an SLM.
In this paper, we propose FastSLM, a lightweight and efficient SLM for processing speech input. To achieve this, we introduce the Hierarchical Frame Querying Transformer (HFQ-Former), a novel module that compresses frame-level features extracted from a speech encoder into compact features. Specifically, HFQ-Former reduces the input length to approximately 1.67 tokens per second by processing 3,000-frame speech chunks. We also propose a three-stage training strategy to jointly optimize performance across speech-based multitasks, including automatic speech recognition (ASR), automatic speech translation (AST), SSUM, and SQQA. Experimental results show that FastSLM attains competitive or even superior performance to state-of-the-art (SOTA) models across multiple benchmarks, despite being trained on comparatively small-scale datasets and with substantially fewer FLOPs. Additionally, to the best of our knowledge, this work introduces the first open-source SLM that supports Korean alongside English, thereby broadening accessibility and applicability to bilingual environments.
The following is a summary of our main contribution:
  • We propose FastSLM, a low-cost and high-efficiency SLM that enables fast and effective text generation from long-form speech inputs with low computational overhead.
  • We introduce the HFQ-Former module, which efficiently compresses frame-level features extracted from a pre-trained speech encoder.
  • We present a three-stage training strategy that effectively adapts a pre-trained LLM for the speech modality, achieving strong alignment between speech and text representations without costly end-to-end training.
  • Through extensive experiments on speech-based multitask evaluation, we demonstrate that FastSLM achieves competitive performance compared to existing SLM, while significantly reducing memory usage.
  • We release the first open-source bilingual (Korean, English) multi-task SLM, providing a powerful and accessible foundation for future research in multilingual speech-language understanding.
@@ 2.1 Audio and Speech Language Models
Although LLMs have achieved human-expert capabilities, their generation is inherently grounded in textual input. Consequently, to advance toward AGI, numerous studies have explored extending LLMs to the audio (including speech) modality. Models such as AudioPaLM (Rubenstein et al., 2023), Kimi-Audio (Ding et al., 2025), Qwen-Audio (Chu et al., 2023), and Qwen2-Audio (Chu et al., 2024) demonstrate that frame-level speech features extracted from a speech encoder can be integrated into the LLM embedding space for end-to-end spoken understanding.
@@ 3.1 Model Architecture of FastSLM
FastSLM: The overall architecture of the proposed FastSLM is illustrated in Fig. 1. FastSLM takes both speech and a text prompt as input to perform text generation. The processing pipeline is as follows. The raw waveform is first converted into a Mel spectrogram using a 25-ms window and a 10-ms stride. The resulting Mel spectrogram is then processed by a speech encoder (Radford et al., 2023) to extract frame-level features at a rate of 50 tokens per second. To use these frame-level features as input to the LLM, they are compressed and temporally aligned into 1.67 tokens per second using the proposed HFQ-Former. Finally, the speech tokens produced by the HFQ-Former are concatenated with the text tokens and provided as input to the LLM. Using this multimodal input, the model generates the final textual response.
@@
Existing Q-Former (Li et al., 2023) is used to generate compact speech representations for the LLM by interacting with features $\mathbf{\hat{X}}^{A}$ extracted by the encoder from the input Mel spectrogram $\mathbf{X}^{A}$ (Tang et al., 2024). In this setup, learnable queries are refined through cross-attention with frame-level features extracted by the speech encoder (Yu et al., 2024). However, relying solely on the speech encoder output $\mathbf{\hat{X}}^{A}$ to learn queries limits the ability to capture the overall context of long-form speech. Furthermore, although segmenting speech features into windowed segments and processing them repeatedly helps to capture local information in $\mathbf{Q}^{A}$, it leads to a rapid increase in computational cost. To address these issues, we propose the HFQ-Former. The structure of HFQ-Former is illustrated in Fig. 2.
@@
HFQ-Former: HFQ-Former adopts a three-stage hierarchical architecture (short-term, mid-term, long-term) to effectively align the high-frame speech features $\hat{\mathbf{X}}^{A}$ with the LLM. In Stage 1, the high-frame speech features $\hat{\mathbf{X}}_{1}^{A}$ retain rich local information and fused with the learnable query $\mathbf{Q}_{1}^{A}$ via cross-attention, producing the intermediate features $\hat{\mathbf{Q}}$. The learnable query $\mathbf{Q}_{i}^{A}$ is used as the query vector, while the speech feature $\hat{\mathbf{X}}_{i}^{A}$ serves as both the key and value. To preserve the positional information of the speech during compression, we incorporate positional encoding $PE(\cdot)$ (He et al., 2022) into the query and key features (Zhang et al., 2025). The cross-attention between $\mathbf{Q}_{i}^{A}$ and $\hat{\mathbf{X}}_{i}^{A}$ is then performed as follows:
@@
$\text{Compressor}(\mathbf{Q}_{i}^{A},\hat{\mathbf{X}}_{i}^{A},\hat{\mathbf{X}}_{i}^{A})=\text{Softmax}\left(\frac{(\mathbf{Q}_{i}^{A}+PE(\mathbf{Q}_{i}^{A}))\cdot(\hat{\mathbf{X}}_{i}^{A}+PE(\hat{\mathbf{X}}_{i}^{A}))^{T}}{\sqrt{d}}\right)\cdot\hat{\mathbf{X}}_{i}^{A},$ (1)
where $d$ denotes the attention dimensionality. In Stage 2 and Stage 3, We perform cross-attention between the learnable query and the speech features. To capture broader temporal context in the speech, we down-sample the high-frame speech features using a Downsampler. Each Downsampler consists of two convolutional layers with kernel size 3 and a GELU activation function (Hendrycks and Gimpel, 2016), where the second convolution uses a stride of 2. This design enables the model to effectively capture both local and global temporal information from speech features.
At each $i$-th stage, the learnable query $\mathbf{Q}_{i}^{A}$ attends to the down-sampled speech features $\hat{\mathbf{X}}_{i}^{A}$ via cross-attention, enabling the model to integrate features at multiple temporal resolutions. The features at each stage can be written as follows:
@@
$\hat{\mathbf{Q}}_{i}^{A}=\text{Compressor}_{i}(\mathbf{Q}_{i}^{A},\hat{\mathbf{X}}_{i}^{A},\hat{\mathbf{X}}_{i}^{A}),~\ \text{where}~\ \hat{\mathbf{X}}_{i}^{A}=\text{Downsampler}_{i}(\hat{\mathbf{X}}_{i-1}^{A}),$ (2)
where $i\in\{1,2,3\}$ and $\hat{\mathbf{X}}_{0}^{A}$ is the extracted high-frame speech features from speech encoder, and C.Attn denotes the cross-attention layer. At each stage, the generated $\hat{\mathbf{Q}}_{i}^{A}$ selectively extracts important information from the speech sequence, and the final representation is obtained by concatenating them as $\hat{\mathbf{Q}}^{A}=[\hat{\mathbf{Q}}_{1}^{A};\hat{\mathbf{Q}}_{2}^{A};\hat{\mathbf{Q}}_{3}^{A}]$. However, when considering long-form speech, the context size of $\hat{\mathbf{Q}}^{A}$ can still be computationally prohibitive for direct use as input to the LLM. We draw inspiration from vision-language models like LLaVA-mini (Zhang et al., 2025), which demonstrated that an entire image can be effectively represented by a single token. However, unlike a static image, speech is a temporal stream with complex, overlapping acoustic events, making extreme compression to a single token challenging without significant information loss. Therefore, to balance representational fidelity with computational efficiency, we compress hierarchical information $\hat{\mathbf{Q}}^{A}$ into a small fixed number of learnable queries $\mathbf{Q}_{c}^{A}$. The output learnable query $\tilde{\mathbf{Q}}_{c}^{A}$ is computed as:
$\tilde{\mathbf{Q}}_{c}^{A}=\text{C.Attn}\left({\mathbf{Q}}_{c}^{A},\hat{\mathbf{Q}}^{A},\hat{\mathbf{Q}}^{A}\right),$ (3)
Subsequently, to capture salient features from the compressed speech tokens $\tilde{\mathbf{Q}}_{c}^{A}$ and to reference hierarchical information across diverse frames, we design an attention block that performs self-attention on $\tilde{\mathbf{Q}}_{c}^{A}$ followed by cross-attention with $\hat{\mathbf{X}}^{A}$. This cross-attention step is crucial as explicitly compensates for potential information loss caused by the dramatic compression of speech features. The output of this attention block, denoted as $\hat{\mathbf{Q}}_{c}^{A}$, is computed as:
@@
where, $\hat{\mathbf{X}}^{A}=[\hat{\mathbf{X}}_{0}^{A};\hat{\mathbf{X}}_{1}^{A};\hat{\mathbf{X}}_{2}^{A}]$ denotes the concatenated frame-level speech features extracted at each stage. This produces the final compressed representation $\hat{\mathbf{Q}}_{c}^{A}$, which integrates both local and global context while reducing sequence length, thereby lowering the computational cost of autoregressive decoding in the LLM.
To further validate the effectiveness of the hierarchical compression, we provide a qualitative analysis of the cross-attention patterns across different stages in Appendix A. The visualization shows that HFQ-Former progressively shifts its attention toward deeper stages when processing long-form speech, confirming that the hierarchical design is essential for long-range temporal abstraction. To further validate the effectiveness of the hierarchical compression, we provide a qualitative analysis of the cross-attention patterns across different stages in Appendix B.
@@ 3.2 Three-Stage Training Strategy
To train FastSLM, we propose a three-stage speech modality adaptation strategy, designed to progressively enhance the model capability to understand and adapt to speech input in LLM. Across all stages, we adopt low-rank adaptation (LoRA) (Hu et al., 2022) to ensure cost-efficient training with minimal trainable parameters. Specifically, we set the LoRA hyperparameter to a rank of 16 and an alpha of 64, resulting in a scaling factor of 4.
@@
Pre-training (Short-form speech Adaptation): In the first stage, the model is trained to adaptation short-form speech inputs. We construct a dataset of approximately 15K hours speech-text pairs in both Korean and English, with each speech clip restricted to under 30 seconds. This ensures that the model can learn general ASR capabilities and effectively align speech with language. We adopt prompt formats inspired by hierarchical tags (Chu et al., 2023; 2024) to improve language-specific understanding and detection during this stage. A detailed description of this can be found in Appendix C.
@@
Long-form Speech Adaptation: Pre-trained speech encoders (Gulati et al., 2020; Radford et al., 2023; Elizalde et al., 2023; Chen et al., 2023) are typically limited to processing speech segments shorter than 30 seconds, which hinders their performance on long-speech tasks such as SSUM, SQQA. To address this limitation, Audio-Flamingo3 (Goel et al., 2025) constructed instruction tuning datasets specifically designed for long-form speech and audio understanding. However, building such datasets requires significant time and cost.
@@
To provide a more cost-effective alternative, we train the model on a curated ASR-based dataset containing speech–text pairs of 1 to 15 minutes in length. This stage is designed not to training abstract reasoning directly but to strengthen the model’s fundamental ability to process extended speech sequences. Through long-form transcription training, the model learns to maintain temporal coherence and preserve acoustic features over lengthy contexts—a critical prerequisite for complex downstream tasks. The resulting long-context representations supply the language model backbone with higher-quality, more coherent inputs, enabling superior performance on tasks such as SSUM and SQQA that require accurate understanding of an entire speech stream. Because ASR datasets are far more accessible than bespoke instruction-tuning dataset, this approach provides a practical and scalable path toward long-form speech modeling.
@@
Instruction Tuning: In the final stage, we perform instruction tuning to enable the model to handle a variety of downstream tasks. Due to the scarcity of multi-task speech-language datasets in non-English languages, we generate a Korean multi-task dataset using the text-to-speech (TTS) engine (Zhao et al., 2023a), covering a range of tasks including SSUM, and SQQA. Unlike the previous stages, hierarchical tags are no longer required, as language identification capabilities have already been sufficiently established. However, hierarchical task tags are still employed to explicitly specify the task. A detailed description of this can be found in Appendix C.
Through this three-stage adaptation process, FastSLM is trained to achieve balanced capabilities across speech modality adaptation, long-form speech comprehension, and multi-task speech language understanding.

### v2

@@ Abstract
Although Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in vision, language, and video understanding tasks, scaling them to long-form speech remains a critical bottleneck due to the explosive growth of input tokens. Existing speech-language models typically project high-frame-rate acoustic features directly into the LLM input space, rendering long-context processing computationally prohibitive as audio duration increases. In this paper, we present FastSLM, a token-efficient architecture designed to overcome this scalability limit through extreme temporal compression. At its core is the Hierarchical Frame Querying Transformer (HFQ-Former), which progressively distills local acoustic details into compact, semantically rich representations across multiple temporal scales. This hierarchical abstraction reduces the speech representation rate to just 1.67 tokens per second—achieving a 93% token reduction compared to standard frame-level adapters—while preserving the critical context required for complex reasoning. Experimental results demonstrate that FastSLM achieves competitive performance with state-of-the-art (SOTA) models on long-form benchmarks, despite operating with significantly lower FLOPs and parameter counts. Our findings establish that extreme token compression is a viable pathway to making real-time, long-context speech understanding feasible for LLMs, even under strict computational constraints. The source code and model checkpoints are available at https://anonymous.4open.science/r/FastSLM-8BD3.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; Comanici et al., 2025; Yang et al., 2025) have demonstrated remarkable reasoning capabilities, prompting a surge in research toward Multimodal LLMs (MLLMs) that perceive and understand the world through vision, audio, and video (Yin et al., 2024; Lyu et al., 2023). Among these modalities, speech is arguably the most primary and natural interface for human-AI interaction (Sakshi et al., 2024). Consequently, adapting LLMs to the speech domain has become a central focus, leading to the development of various Audio Language Models (ALMs) (Tang et al., 2024; Abouelenin et al., 2025; Chu et al., 2023, 2024; Goel et al., 2025).
However, a fundamental challenge hinders the scalability of current ALMs: the granularity mismatch between acoustic features and textual tokens. Prevalent approaches typically project frame-level features from encoders, such as Whisper (Radford et al., 2023) and Conformer (Gulati et al., 2020), directly into the LLM input space. While effective for short clips, this dense projection strategy generates hundreds of tokens for just a few seconds of audio. As the input duration extends to minutes or hours, the sequence length explodes, causing the computational cost of the autoregressive LLM, which scales quadratically ($O(N^{2})$), to become prohibitive. This phenomenon, often referred to as the token explosion, leads to excessive KV cache growth and unacceptable latency, rendering real-time processing impractical (Arif et al., 2025).
This bottleneck is particularly critical given the growing demand for long-form speech understanding in real-world applications such as meeting summarization (speech summarization (SSUM)) and spoken query-based question answering (SQQA) (Ghosh et al., 2024; Kang and Roy, 2024). Unlike short commands, these tasks require the model to reason over long temporal contexts containing thousands of acoustic events. Existing architectures, constrained by their dense token representation, struggle to handle such long sequences efficiently. To address this, we propose a paradigm shift from simple feature alignment to extreme temporal compression.
To bridge this gap, we present FastSLM, a cost-effective architecture designed to distill long-form speech into highly compact representations, making long-context reasoning feasible even under strict computational constraints. To drive this architecture, we introduce the Hierarchical Frame Querying Transformer (HFQ-Former), a novel mechanism that departs from flat projection by hierarchically compressing acoustic features. By progressively abstracting local details into global contexts, HFQ-Former reduces the speech representation rate to approximately 1.67 tokens per second, enabling the processing of hour-long audio on consumer-grade hardware. Furthermore, we introduce a three-stage training strategy designed to maximize data efficiency, enabling robust long-context adaptation without relying on expensive human-annotated long-form instruction datasets.
Our main contributions are summarized as follows:
  • We propose FastSLM, a streamlined architecture that solves the “token explosion” problem in MLLMs, enabling real-time reasoning over long-form speech with minimal computational overhead.
  • We introduce the HFQ-Former, which achieves extreme token compression (approximately 93% reduction vs. frame-level features) via hierarchical cross-attention, effectively balancing fine-grained acoustic perception with global semantic understanding.
  • We demonstrate a cost-effective three-stage training strategy that aligns pre-trained LLMs with long-form speech using only standard ASR corpora, bypassing the need for scarce long-context instruction data.
  • Empirical results show that FastSLM achieves competitive or superior performance to SOTA models on benchmarks such as SSUM and SQQA while exhibiting near-linear memory scaling.
@@ 2.1 Audio and Speech Language Models
Although LLMs have achieved human-expert capabilities, their generation is inherently grounded in textual input. Consequently, to advance toward artificial general intelligence (AGI), numerous studies have explored extending LLMs to the audio (including speech) modality. Models such as AudioPaLM (Rubenstein et al., 2023), Kimi-Audio (Ding et al., 2025), Qwen-Audio (Chu et al., 2023), and Qwen2-Audio (Chu et al., 2024) demonstrate that frame-level speech features extracted from a speech encoder can be integrated into the LLM embedding space for end-to-end spoken understanding.
@@ 2.2 Q-Former for Speech Modality Adaptation
Existing Q-Former approaches (Li et al., 2023; Tang et al., 2024) typically utilize learnable queries to extract static features from fixed, short-term windows or simple segments (Yu et al., 2024). However, this design suffers from a context-granularity trade-off: applying Q-Formers to short windows risks missing global context, whereas applying them to long segments tends to dilute local acoustic details. Furthermore, naive segmentation leads to a linear increase in token count, failing to address the scalability challenges inherent in long-form speech. To resolve these issues, we propose the HFQ-Former, which introduces a progressive abstraction mechanism to efficiently capture both local details and global context.
@@ 3.1 Model Architecture of FastSLM
FastSLM Overview: The overall architecture of FastSLM is illustrated in Fig. 1. Unlike conventional SLMs that directly concatenate dense acoustic features with text embeddings, FastSLM introduces a dedicated compression bottleneck to bridge the information density gap between modalities. The pipeline proceeds as follows: Raw waveforms are converted into Mel spectrograms (25-ms window and a 10-ms stride) and processed by a speech encoder (Radford et al., 2023) to extract high-fidelity frame-level features (50 Hz). Crucially, instead of feeding these dense frames to the LLM, we employ the HFQ-Former to distill them into highly compact semantic tokens. This mechanism achieves an extreme compression rate of approximately 1.67 tokens per second, reducing the sequence length by over 96% compared to the original encoder output. These compressed tokens are then concatenated with text prompts, allowing the LLM to perform long-context reasoning with minimal computational overhead.
@@
HFQ-Former (Hierarchical Frame Querying Transformer): The HFQ-Former adopts a three-stage hierarchical design to progressively abstract speech information from acoustic to semantic levels. This mimics the cognitive process of progressively summarizing continuous acoustic signals into compact semantic representations. The detailed structure of HFQ-Former is shown in Fig. 2.
Stage 1 (Local Acoustic Extraction): The first stage focuses on preserving fine-grained local information. The high-frame-rate features $\hat{\mathbf{X}}_{1}^{A}\in\mathbb{R}^{F}$ from the encoder are fused with a set of learnable queries $\mathbf{Q}_{1}^{A}$ via cross-attention. Here, $\mathbf{Q}_{1}^{A}$ acts as a local probe, extracting phoneme-level details while maintaining the original temporal resolution. To preserve order during this dense interaction, we inject sinusoidal positional encoding $PE(\cdot)$ (He et al., 2022) into both queries and keys. The compression operation at stage $i$ is defined as:
@@
$\text{Compressor}(\mathbf{Q}_{i}^{A},\hat{\mathbf{X}}_{i}^{A},\hat{\mathbf{X}}_{i}^{A})$ $=\text{Softmax}\!\left(\frac{\tilde{\mathbf{Q}}_{i}^{A}\tilde{\mathbf{X}}_{i}^{A\!T}}{\sqrt{d}}\right)$ (1)
$\cdot\hat{\mathbf{X}}_{i}^{A},\text{where }\tilde{\mathbf{Q}}_{i}^{A}$ $=\mathbf{Q}_{i}^{A}+PE(\mathbf{Q}_{i}^{A}),$
$\tilde{\mathbf{X}}_{i}^{A}$ $=\hat{\mathbf{X}}_{i}^{A}+PE(\hat{\mathbf{X}}_{i}^{A}),$
Stages 2 & 3 (Temporal Aggregation & Semantic Abstraction): Stage 1 captures local acoustic details, while Stages 2 and 3 progressively abstract this information into higher-level semantic representations. A Downsampler acts as a temporal filter, reducing redundancy (e.g., silence, stationary noise). Each Downsampler consists of dual convolutional layers with kernel size 3 and GELU activation (Hendrycks and Gimpel, 2016), applying a stride of 2 to halve temporal resolution at each step. Learnable queries $\mathbf{Q}_{i}^{A}$ at deeper stages attend to broader temporal contexts, transitioning from phoneme-level (Stage 1) to word/phrase-level (Stage 2), and finally to sentence-level semantics (Stage 3). The hierarchical feature update is formalized as:
@@
$\begin{split}\hat{\mathbf{Q}}_{i}^{A}=\text{Compressor}_{i}(\mathbf{Q}_{i}^{A},\hat{\mathbf{X}}_{i}^{A},\hat{\mathbf{X}}_{i}^{A}),\\ \text{where}\quad\hat{\mathbf{X}}_{i}^{A}=\text{Downsampler}_{i}(\hat{\mathbf{X}}_{i-1}^{A}).\end{split}$ (2)
where $i\in\{2,3\}$, and $\hat{\mathbf{X}}_{0}^{A}$ denotes the initial encoder output. Multi-scale representations are concatenated as $\hat{\mathbf{Q}}^{A}=[\hat{\mathbf{Q}}_{1}^{A};\hat{\mathbf{Q}}_{2}^{A};\hat{\mathbf{Q}}_{3}^{A}]$ to form a comprehensive feature bank.
Semantic Distillation (Extreme Token Compression): Although $\hat{\mathbf{Q}}^{A}$ contains rich information, its length scales linearly with audio duration, which remains suboptimal for long-form reasoning. To achieve our target efficiency of 1.67 tokens/sec, we introduce a final Semantic Distillation step. Drawing inspiration from LLaVA-mini (Zhang et al., 2025), which compresses visual features into a single token, we hypothesize that speech segments can also be represented by a highly compact set of latent variables. However, unlike static images, speech represents a dynamic temporal stream with overlapping events. Therefore, instead of a single token, we employ a small, fixed number of learnable queries $\mathbf{Q}_{c}^{A}$ to distill the hierarchical information $\hat{\mathbf{Q}}^{A}$ into the final compressed tokens $\tilde{\mathbf{Q}}_{c}^{A}$:
$\tilde{\mathbf{Q}}_{c}^{A}=\text{C.Attn}\left({\mathbf{Q}}_{c}^{A},\hat{\mathbf{Q}}^{A},\hat{\mathbf{Q}}^{A}\right).$ (3)
Detail Recovery Mechanism: Extreme compression risks losing subtle but critical acoustic cues (e.g., intonation, speaker identity). To mitigate this, we incorporate a Detail Recovery attention block. This module allows the compressed semantic tokens $\tilde{\mathbf{Q}}_{c}^{A}$ to look back at the raw multi-scale source features $\hat{\mathbf{X}}^{A}$ and retrieve any missing fine-grained details before entering the LLM.
@@
where $\hat{\mathbf{X}}^{A}=[\hat{\mathbf{X}}_{0}^{A};\hat{\mathbf{X}}_{1}^{A};\hat{\mathbf{X}}_{2}^{A}]$. The final output $\hat{\mathbf{Q}}_{c}^{A}$ provides a strong balance between semantic density and acoustic fidelity, addressing the granularity mismatch described in Section 1.
To substantiate the efficacy of our hierarchical compression mechanism, we provide a qualitative analysis of cross-attention patterns in Appendix A. The visualizations demonstrate that HFQ-Former progressively shifts its attention toward deeper stages as the speech duration increases, confirming that the proposed hierarchical architecture is essential for effective long-range temporal abstraction. Furthermore, we present additional analyses on hierarchical temporal modeling in Appendix B and investigate the impact of the intermediate query size ($\hat{\mathbf{Q}}_{i}^{A}$) in Appendix C.
@@ 3.2 Three-Stage Training Strategy
To train FastSLM, we propose a three-stage speech modality adaptation strategy [Image of three-stage training pipeline for speech language models showing pre-training, long-form adaptation, and instruction tuning], designed to progressively enhance the model’s capability to understand and adapt to speech input. Across all stages, we adopt low-rank adaptation (LoRA) (Hu et al., 2022) to ensure cost-efficient training with minimal trainable parameters. Specifically, we set the LoRA rank to 16 and alpha to 64, resulting in a scaling factor of 4.
@@
Pre-training (Short-form Speech Adaptation): In the first stage, the model is trained to adapt to short-form speech inputs. We construct a dataset of approximately 17K hours of speech-text pairs in both Korean and English, with each speech clip restricted to under 30 seconds. This ensures that the model can learn general ASR capabilities and effectively align speech with language. We adopt prompt formats inspired by hierarchical tags (Chu et al., 2023, 2024) to improve language-specific understanding during this stage. A detailed description can be found in Appendix D.
@@
Long-form Speech Adaptation: Pre-trained speech encoders (Gulati et al., 2020; Radford et al., 2023; Elizalde et al., 2023; Chen et al., 2023) are typically limited to processing segments shorter than 30 seconds, hindering performance on long-speech tasks such as SSUM and SQQA. While Audio-Flamingo3 (Goel et al., 2025) addressed this by constructing bespoke instruction tuning datasets, such an approach is time-consuming and costly.
@@
To provide a cost-effective alternative, we train the model on a curated ASR-based dataset containing speech–text pairs ranging from 1 to 15 minutes. This stage is designed not to train abstract reasoning directly but to strengthen the model’s fundamental ability to process extended speech sequences. Through long-form transcription training, the model learns to maintain temporal coherence and preserve acoustic features over lengthy contexts—a critical prerequisite for complex downstream tasks. The resulting long-context representations supply the LLM backbone with higher-quality, more coherent inputs, enabling superior performance on downstream tasks. Since ASR datasets are far more accessible than specialized instruction-tuning datasets, this approach offers a scalable path toward long-form speech modeling.
@@
Instruction Tuning: In the final stage, we perform instruction tuning to enable the model to handle a variety of downstream tasks. Due to the scarcity of non-English multi-task datasets, we generated a Korean multi-task dataset using a text-to-speech (TTS) engine (Zhao et al., 2023), covering tasks such as SSUM and SQQA. Unlike previous stages, hierarchical language tags are no longer required as language identification is established; however, hierarchical task tags are employed to explicitly specify the task. Details are provided in Appendix D.